OpenStudy (anonymous):

If x2 + y2 = 14 and xy = 3, then (x + y)2 =?

OpenStudy (anonymous):

Try expanding this:\[\Huge(x+y)^2=\cdots\]

OpenStudy (anonymous):

\[(x + y)^{2} = x ^{2} + y ^{2} + 2xy\] Now just plug values in. \[x ^{2} + y ^{2} + 2xy = 14 + 2(3) = 20\]

OpenStudy (anonymous):

(x+y)^2=x^2+y^2+2xy Substitute x^2+y^2=14 and xy=3 (x+y)^2=14+2(3)=20